### v1

# AGRAG: Advanced Graph-based Retrieval-Augmented Generation for LLMs

Abstract
Graph-based retrieval-augmented generation (Graph-based RAG) has demonstrated significant potential in enhancing Large Language Models (LLMs) with structured knowledge. However, existing methods face three critical challenges: Inaccurate Graph Construction, caused by LLM hallucination; Poor Reasoning Ability, caused by failing to generate explicit reasons telling LLM why certain chunks were selected; and Inadequate Answering, which only partially answers the query due to the inadequate LLM reasoning, making their performance lag behind NaiveRAG on certain tasks. To address these issues, we propose AGRAG, an advanced graph-based retrieval-augmented generation framework. When constructing the graph, AGRAG substitutes the widely used LLM entity extraction method with a statistics-based method, avoiding hallucination and error propagation. When retrieval, AGRAG formulates the graph reasoning procedure as the Minimum Cost Maximum Influence (MCMI) subgraph generation problem, where we try to include more nodes with high influence score, but with less involving edge cost, to make the generated reasoning paths more comprehensive. We prove this problem to be NP-hard, and propose a greedy algorithm to solve it. The MCMI subgraph generated can serve as explicit reasoning paths to tell LLM why certain chunks were retrieved, thereby making the LLM better focus on the query-related part contents of the chunks, reducing the impact of noise, and improving AGRAG’s reasoning ability. Furthermore, compared with the simple tree-structured reasoning paths, our MCMI subgraph can allow more complex graph structures, such as cycles, and improve the comprehensiveness of the generated reasoning paths.
I Introduction
Figure 1: An overview of Graph-based RAG models. It first extracts entities and relations as graph nodes and edges. When retrieval, they will map query to graph nodes, and utilizing the graph structure to support query answering.
Large Language Models (LLMs) [1, 2, 3, 4, 5] have demonstrated remarkable capabilities in natural language understanding and generation [6, 7, 8, 9, 10, 11, 12, 13, 14]. However, their reliance on static, internalized knowledge often leads to factual hallucinations and limited adaptability to dynamic information [15, 16, 17, 18], and retraining or finetuning the LLM can lead to high computational costs [19, 20]. Retrieval-Augmented Generation (RAG) addresses this by integrating external factual knowledge during inference, improving accuracy and enabling real-time updates. Yet, traditional RAG methods’ retrieval [21, 15] is based on simplistic mechanisms, such as the whole passages’ dense vector similarity, make them overlook the capture of more fine-grained semantic relationships between entities [22, 23, 24]. To overcome these limitations, researchers have proposed various Graph-based RAG models [25, 26, 27, 28, 29, 30, 31]. The framework of Graph-based RAG models, is shown in Figure 1. These models first utilize LLMs to extract entities and relations from the source documents to serve as graph nodes and edges and construct a graph, and then leverage this graph-structured knowledge to model complex relationships between entities to enable structured reasoning over inter-connected information, enabling the capture of more fine-grained entity-level relationships when retrieval. [22, 24].
Currently, the Graph-based RAG models can be classified into two categories, i.e., community-based models, and walking-based models. The community-based models [25, 28, 26] first map the query to graph nodes considering their semantic similarity or keyword matching. Then, they will retrieve the mapped graph nodes’ one or multi-hop neighborhood. On the other hand, with similar mapping approach from query to graph nodes, walking-based models [29, 32, 27, 30, 31] apply walk-based graph algorithms, such as Personalized PageRank (PPR) [33], based on the mapped graph nodes, to score graph nodes and edges based on the graph’s topological structure and query information, and then retrieve Top-k text chunks based on high-scoring graph nodes, capturing the fine-grained entity level relationships.
However, on complex QA tasks that require knowledge summarization or content generation [34], the existing RAG models cannot answer these complex queries due to:
  1. 1.
    Inaccurate Graph Construction: Current Graph-based RAG models apply LLM for entity and relation extraction to construct graphs. However, the LLM can inevitably suffer from hallucination issue [15, 16, 17, 18] and leading to inaccurate entity extraction, introducing noise into the constructed graph and propagate through the whole pipeline, affecting the model effectiveness.
  2. 2.
    Poor Reasoning Ability: The community-based models’ retrieval only focuses on a narrow graph neighborhood and misses the graph’s long-hop dependency, which is crucial for tasks that require multi-hop reasoning. In contrast, walking-based models do account for these long-hop dependencies during retrieval; however, they simply feed the retrieved text chunks directly into the LLM as contexts. Consequently, both of their retrieval and reasoning processes remain opaque to the LLM, which receives no explicit structural reasoning chain to explain why those particular chunks were selected, making the LLM struggle to locate the query-related context within those chunks, leading to suboptimal query answering performance on certain tasks [34, 35, 36].
  3. 3.
    Inadequate Answer: Although we can simply combine the existing RAG models with the current shortest path-based graph algorithm to provide explicit reasoning chains to aid LLM query answering, improve the model’s reasoning ability. However, on tasks that require summarization of current knowledge [34, 35, 36], this simple approach can lead to suboptimal performance, as the shortest path on the graph can be overly simple, and can omit many query-related facts, leading to incomplete reasoning chains, which makes the model tend to generate inadequate query answer with low information coverage and faithfulness (See Table VI and subsection IV-D).
To address these issues, we propose a novel Advanced Graph-based Retrieval-Augmented Generation framework, called AGRAG. We divide the AGRAG framework into three steps: In Step 1 Data Preparation, AGRAG first extracts entities with a TFIDF [37] based entity extraction method, avoiding LLM hallucination, and resulting in better graph construction accuracy. Then, AGRAG applies LLM to detect the possible relation w.r.t. the extracted entity and the text chunk, formulating entity-relation-entity triples, namely knowledge triple. Then these knowledge triples, as well as the text chunks involved, together formulate a Knowledge Graph (KG). In Step 2 Graph Retrieval, AGRAG maps the query to knowledge triples in KG based on their semantic similariy, then calculate the PPR score [33] as node influence based on the KG structure and the mapped triples. In the meantime, AGRAG also calculates edge weights based on the semantic similarity between the query and the edge’s corresponding triple. After that, AGRAG generates a Minimum Cost and Maximum Influence (MCMI) reasoning subgraph via. a greedy algorithm, which is calculated based on the shortest spanning tree of the mapped triples. It tries to maximize the node influence score, while minimize the edge cost in the expansion of each node, in order to avoid introducing noise. Compared to other shortest path-based algorithms, the MCMI subgraph allows for more complex structures, such as rings, which results in reasoning chains containing more query-related entities and relations, making them more comprehensive. This, in turn, helps the model perform better on complex tasks that require summarization, resulting in better reasoning ability and more comprehensive query answers. In Step 3 Hybrid Retrieval, AGRAG carries out hybrid retrieval based on the average of dense representation vector similarity score and sparse BM25 [38] keyword matching score between query and text chunks, providing extra contexts for query answering. At last, AGRAG will input the resulting subgraph from the Graph Retrieval step and chunks from hybrid retrieval to LLM for final query answering.
We summarize our novel contributions as follows:
  • We present the AGRAG framework, which consists of three steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval, aiming at addressing the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current Graph-based RAG models.
  • We design a statistic-based entity extraction method, avoiding the hallucination caused by LLM entity extraction, and improving the pipeline effectiveness.
  • We propose the MCMI subgraph generation problem, prove it to be NP-hard, and design a greedy algorithm, ensuring AGRAG’s generalizability to complex tasks.
  • According to our experiments, AGRAG achieve better effectiveness and efficiency compared with current state-of-the-art RAG models on six tested tasks.
II Related Works
In this section, we first introduce the traditional RAG models in Section II-A and then discuss the more advanced Graph-based RAG models in Section II-B. The important notations used in this paper are listed in Table I.
II-A Traditional RAG Models
Recently, Large Language Model (LLM)-based models have undergone rapid development [39, 40, 41, 42, 43, 44, 45, 46, 47] and have been successfully adapted to various data base related tasks, including data discovery [48, 49, 50], entity or schema matching [51, 52, 53, 54], and natural language to SQL conversion [55, 56, 57, 58].
Although LLMs are inherently capable of inference without fine-tuning [1, 2, 3, 4, 5]. However, the lack of fine-tuning with up-to-dated and domain specific knowledge can lead LLMs to generate incorrect answers, as they lack task-specific knowledge, these incorrect answers are often referred to as hallucinations [59]. To mitigate hallucinations, researchers try to directly retrieve text documents outside of LLM based on the query as side information to help the LLM query answering [60, 61, 62, 63, 64, 65, 66, 67], lead to the emergence of early RAG models, namely traditional RAG models. The earliest LLM-RAG model is the simple NaiveRAG [15], when indexing, it split the texts into chunks, and encode them to semantic vector representations. When retrieval, the model retrieve several text chunks with top semantic similarity to the query.
Based on NaiveRAG, researchers have walking-based various models to further tailor the text chunks and improve the retrieval quality. For example, Propositionizer [60] applies a fine-tuned LLM to convert the text chunks into atomic expressions, namely propositions, to facilitate fine-grained information retrieval. Prompt compressor models [68, 69], such as LongLLMLingua [68] apply LLM’s generation perplexity on to filter out un-important tokens in the text chunks, try to maintain a more concise retrieval result for LLM input. However, the retrieval procedures of these models remain unstructured and struggle to capture the nuanced semantic and structured relationship within the text chunks, which impedes the effectiveness of LLMs [15, 16, 17, 18].
II-B Graph-based RAG Models
To overcome the aforementioned limitations of traditional RAG models, researchers have begun to develop Graph-based RAG models [25, 26, 27, 28, 29, 30, 31]. These models usually extract graph components with LLMs from text corpis to index them into graphs. When retrieval, these models will utilize the graph topology structure to capture complex relationships between entities within and across text chunks.
II-B1 Community-based Models
Early Graph-based RAG models typically retrieve specific neighborhoods from the graph based on the input query. The pioneering work in this category is GraphRAG [25], as illustrated in Figure 1, which constructs a graph and aggregates nodes into communities, followed by generating community summaries to assist the LLM in making predictions. However, GraphRAG requires re-computing the graph communities after each index update, leading to significant computational overhead.
To address this limitation, LightRAG [26] eliminates the explicit detection of graph communities, as well as the computation of community summaries. Instead, it retrieves the most semantically similar entities along with their 1-hop graph neighborhoods and the corresponding text passages containing these entities as external knowledge, which are then fed to the LLM. Nevertheless, this approach can still incur computational costs and may lead to hallucinations, as the 1-hop neighborhoods can be large and often contain noisy information.
Moreover, all these community-based models are limited in their ability to capture structural information beyond a small local neighborhood, thereby restricting their ability for long-range reasoning.
TABLE I: Summary of important notations.
Notations Meanings
$t$ The text chunk.
$q$ The input query.
$\mathcal{T}$, $\mathcal{T}_{c}$ Original texts and splited text chunks.
$l_{t}$, $l_{o}$ Text chunk and token overlap length.
$s$, $c$ The node influence score and edge cost.
$\mathcal{S}_{\mathcal{V}}$, $\mathcal{C}_{\mathcal{E}}$ Set of node influence score and edge cost.
$\tau$, $\phi$ Entity extraction and synonym threshold.
$b$ The maximum n-gram in extity extraction.
$k_{a}$, $k_{r}$ Number of raw mapped triples, and hybrid retrieved chunks.
$f$ The mapped triplet fact in KG.
$\mathcal{F}_{\mathcal{KG}}$ All triplet facts within the KG.
$\mathcal{F}_{raw}$, $\mathcal{F}$ Raw and mapped triplet facts after filter.
$v$ A node (entity) in KG.
$e_{(u,v)}$ An edge (relation) between node $u$ and $v$.
$\mathcal{KG}(\mathcal{V},\mathcal{E})$ KG with entity set $\mathcal{V}$ and relation set $\mathcal{E}.$
$d$, $\mathbf{p}$, $\mathbf{P}$ The damping factor, personalization vector and transition probablity matrix of PPR.
$\mathcal{N}(\cdot)$ 1-hop neighbor set of a node in KG.
$r_{MCMI}$ MCMI subgraph’s average score-cost ratio.
$g$ MCMI subgraph’s text-formed graph string.
$p_{re}$, $p_{tf}$, $p_{gen}$ LLM prompt for relation extraction, triplet filtering, and answer generation.
$AS(\cdot)$, $HS(\cdot)$ The anchor score and hybrid retrieval score.
$EC(\cdot)$ The edge cost function.
II-B2 Walking-based Models
To further enhance the long-range graph reasoning capabilities of community-based models, researchers have proposed walking-based models [29, 32, 27, 31], which apply walking-based graph algorithms, to detect and capture global topological information within the Knowledge Graphs (KGs).
Among these, HippoRAG [29] and Fast-GraphRAG [32] utilize the Personalized PageRank (PPR) algorithm to traverse the constructed KG and assign relevance scores to graph nodes with respect to the input query. During retrieval, these models select text chunks based on the aggregated PPR scores of the graph nodes they contain. StructRAG [27] trains a router module that learns to identify the optimal reasoning path over the graph based on the input query. HippoRAG2 [31] further improves upon HippoRAG by incorporating text chunks as passage nodes directly into the KG. It retrieves text chunks based on the individual PPR scores of their corresponding passage nodes within the graph.
In this paper, we propose a novel graph-based and efficient RAG framework, termed AGRAG. AGRAG first constructs a KG using a statistics-based entity extraction method, and then assigns graph’s node weights based on both the KG structure and the input query. Finally, it generates a Minimum Cost and Maximum Influence (MCMI) reasoning subgraph to provide the LLM with explicit reasoning chains that maximize query-relevant information while minimizing noise during retrieval.
III Methodology
In this section, we first introduce the overall AGRAG framework, then detail its three core steps: Data Preparation, Graph Retrieval, and Hybrid Retrieval. Finally, we discuss the computational hardness and complexity analysis of AGRAG.
Figure 2: An overview of AGRAG. In Step 1, AGRAG constructs a KG based on the text corpus. In Step 2, the query is mapped to relevant triplet facts in KG, then node score and edge weights are assigned based on PPR algorithm and the semantic similarity between query and each fact, respectively. In Step 3, AGRAG generates a MCMI subgraph, providing explicit reasoning chains in addition the most semantically similar text chunks by Hybrid Retrieval (HR) to support LLM query answering.
III-A Framework Overview
As shown in Figure 2, AGRAG consists of three steps, i.e., Data Preparation, Graph Retrieval, and Hybrid Retrieval.
Step 1: Data Preparation. In this step, AGRAG constructs a Knowledge Graph (KG) $\mathcal{KG}(\mathcal{V},\mathcal{E})$ based on the given text corpus $\mathcal{T}$. Here, we substitute other Graph-based RAG baselines’ LLM entity extraction to a statistics-based method based on TFIDF [37], and extract entities from text chunks in $\mathcal{T}_{c}$ that splitted from the texts in corpus $\mathcal{T}$. This method can avoid LLM hallucination, and hence reduce the noise in the constructed graph, leading to better effectiveness. After entity extraction, we apply LLM to detect relation edges w.r.t. the extracted entity nodes and the text chunk. Furthermore, we also construct text chunks into KG as passage nodes, and link them to all their extracted entity nodes with node-to-passage edges. The detailed pseudo code of Step 1 is shown in algorithm 1, and more details can be found in Section III-B.
Step 2: Graph Retrieval. In this step, we firstly map the query $q$ to KG fact triples considering their semantic similarities, after that, we calculate the PPR score [33] based on these mapped fact triples $F$ as the node influence score $\mathcal{S}_{\mathcal{V}}$. On the other hand, we calculate the edge costs $\mathcal{C}_{\mathcal{E}}$ considering the semantic similarity between query and facts’ vector representations. the higher their similarity is, the lower the cost is. Compared with solely applying PPR to score nodes for retrieval, the edge cost further provides constraints on the involved nodes based on the semantic correspondence between query and facts, removing the need for the human-defined PPR retrieval threshold. After that, we initialize the MCMI subgraph with the generated MCST, and apply our greedy algorithm by iteratively expanding the MCMI subgraph by selecting its neighboring node with the highest influence score-edge cost ratio (i.e., $s/c$). If the highest ratio is smaller than the average ratio of the current MCMI, the iteration stops and the current MCMI subgraph is returned. We prove the MCMI generation problem to be NP-hard, and our greedy algorithm in algorithm 3 can achieve 2-approximate. Please refer to Section III-C for further details.
Step 3: Hybrid Retrieval. In this step, we perform hybrid retrieval and obtain text chunks based on their BM25 sparse similarity score, and dense vector representation similarity score with the query. These chunks will be input together with the MCMI subgraph to the LLM for the final query answering, providing extra contexts.
III-B Step 1: Data Preparation
In this subsection, we utilize a TFIDF [37] based entity extraction approach to substitute the LLM-based approach of other Graph-based RAG models, reduce LLM hallucination, resulting in less noise generated from the early entity extraction step, and hence less error propagation, address the inaccurate graph construction issue, lead to better effectiveness. The pseudo code of AGRAG’s graph construction procedure is shown in algorithm 1.
Input : $\mathcal{T}$: The text corpus for graph construction.
Output : $\mathcal{KG}(\mathcal{V},\mathcal{E})$: The KG constructed.
1 Let $\mathcal{V}\leftarrow\emptyset$, $\mathcal{E}\leftarrow\emptyset$;
2 Get text chunk set $\mathcal{T}_{c}$ from $\mathcal{T}$: $\mathcal{T}_{c}\leftarrow\text{Chunking}(\mathcal{T})$;
3 for each text chunk $t\in\mathcal{T}_{c}$ do
4   Extract entity set $\mathcal{V}_{t}$ based on Equation 3;
5   Carry out relation extraction with LLM: $\mathcal{E}_{t}\leftarrow\text{LLM}(\mathcal{V}_{t},t,p_{\text{re}})$;
6 $\mathcal{V}\leftarrow\mathcal{V}\cup\mathcal{V}_{t}$, $\mathcal{E}\leftarrow\mathcal{E}\cup\mathcal{E}_{t}$;
7
8Add $\mathcal{T}_{c}$ to $\mathcal{V}$ as passage nodes: $\mathcal{V}\leftarrow\mathcal{V}\cup\mathcal{T}_{c}$;
9 for each entity node $v\in\mathcal{V}$ do
10 for each text chunk $t\in\mathcal{T}_{c}$ do
11    if $v\in t$ then
12         Add passage edge between $v$ and $t$: $\mathcal{E}\leftarrow\mathcal{E}\cup\{e_{(v,t)}\}$;
13
14
15
16return $\mathcal{KG}(\mathcal{V},\mathcal{E})$;
Algorithm 1 Indexing Algorithm of AGRAG
To begin with, we split texts in $\mathcal{T}$ into fix lengthed chunks with a certain amount of token overlap between chunks:
$\mathcal{T}_{c}=\text{Split}(\mathcal{T},l_{t},l_{o}),$ (1)
where $l_{t}$ and $l_{o}$ denotes the length of text chunks and token overlap between chunks. Next, we apply the statistics-based entity extraction approach on the text chunks $\mathcal{T}_{c}$. Specifically, our approach modifies the TFIDF score [37] as the entity extraction score $\text{ER}(v,t)$ for each term $v$ in text chunk $t$ as an entity, to make the value of ER fall in $(0,1)$. Here, our entity detection is based on an n-gram setting with the maximum n-gram (number of words in an entity) as $b$, and the equation of $\text{ER}(v,t)$ is as follow:
$\text{ER}(v,t)=\frac{\text{count}(v,t)}{|t|log(|\mathcal{T}_{c}|+1)}\times log\frac{|\mathcal{T}_{c}|+1}{|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}_{c}|+1},$ (2)
where $\mathcal{T}_{c}$ denotes all text chunks obtained from Equation 1, $\text{count}(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}_{c}|$ denotes the number of texts in the corpus $\mathcal{T}_{c}$ that contain the keyword $v$. Here, our entity extraction procedure will extract words or phrases with $\text{ER}(v,t)>\tau$ as the entity nodes $\mathcal{V}_{t}$ for text chunk $t$:
$\mathcal{V}_{t}=\left\{v\,\middle|\,\text{ER}(v,t)>\tau\right\}.$ (3)
Entities in $\mathcal{V}_{t}$ are later served as graph nodes, hence we denote them as entity nodes. For our entity extraction procedure, the entity boundary $b$ and entity extraction threshold $\tau$ are considered as hyper-parameters, we further analyze their setting in subsection IV-E.
After entity extraction, we extract relations from each text chunk $t\in\mathcal{T}_{c}$ based on its extracted entity $\mathcal{V}_{t}$. Similar from other Graph-based RAG models, we also apply LLM at the relation extraction step:
$\mathcal{E}_{t}=\text{LLM}(\mathcal{V}_{t},\ t,\ p_{\text{re}}),$ (4)
where $\mathcal{V}_{t}$ denotes the extracted entities of text chunk $t$, and in the relation extraction prompt $p_{\text{re}}$, we ask the LLM to detect relations between entities in $\mathcal{V}_{t}$ w.r.t. text chunk $t$. Although we also apply LLM for relation extraction, we input less noise in this step, mitigating the error propagation, as our statistics-based entity extraction avoids LLM hallucination, hence resulting in a more comprehensive and less noisy $\mathcal{V}_{t}$ as the Equation 4 input.
After relation extraction, we add text chunks in $\mathcal{T}_{c}$ to $\mathcal{V}$ as passage nodes, and add passage edges $e_{(v,t)}$ between entities the text chunks contain them to $\mathcal{E}$, the respective relation of these edges are namely Contains. By introducing passage nodes, we integerate contextual information into the constructed KG, and the entity-passage edges can capture the inter-text correlations, make our model able to generate reasoning paths across different text chunks [31].
Furthermore, we add synonym edges $\mathcal{E}_{\text{syn}}$ between entity nodes in $\mathcal{V}$ that have similarties larger than the synonym threshold $\phi$, the relation of these edges are namely synonym:
$\mathcal{E}_{\text{syn}}=\left\{e_{(u,v)}\,\middle|\,\text{cosine\_sim}(\textbf{u},\textbf{v})>\phi\right\},$ (5)
At last, we combine edge set $\mathcal{E}_{t}$ for all text chunks $t$ and passage edges together, as well as the synonym edges $\mathcal{E}_{\text{syn}}$, to form the total edge (relation) set of the KG:
$\mathcal{E}=\bigcup_{t\in\mathcal{T}_{c}}\mathcal{E}_{t}\ \cup\bigcup_{v\in\mathcal{V}_{t},\ t\in\mathcal{T}_{c}}e_{(v,t)}\ \cup\mathcal{E}_{\text{syn}},$ (6)
here, if the entity $v$ not exist in text chunk $t$, $e_{(v,t)}=\emptyset$.
Figure 3: An example of the MCMI generated. Red triples indicate those mapped from the query. Our algorithm first constructs the MCMI based on a Steiner tree rooted at these mapped triples, then iteratively expands it by adding neighboring triples with the highest influence-cost ratio $s/c$, continuing until no neighboring triples have this ratio greater than the average influence-cost ratio of the current graph, covering more query related triples (in orange).
III-C Step 2: Graph Retrieval
In this subsection, we introduce the graph retrieval procedure of AGRAG. Given the KG $\mathcal{KG}(\mathcal{V},\mathcal{E})$ constructed at the Step 1, we first map the query $q$ to $\mathcal{KG}$ and obtain raw mapped triplet facts $\mathcal{F}_{\text{raw}}$. Then we encode each triplet fact in KG $f\in\mathcal{F}_{\mathcal{KG}}$ and the query $q$ into vector representations q and f with encoder $\mathcal{M}$: $\textbf{q}\leftarrow\mathcal{M}(q)$, then select facts with top-k dense semantic mapping score MS to query $q$ based on their cosine similarity:
$\text{MS}(\textbf{q},\textbf{f})=\text{cosine\_sim}(\textbf{q},\textbf{f}),$ (7)
$\mathcal{F}_{\text{raw}}=\left\{f\in\mathcal{F}_{\mathcal{KG}}\,\middle|\,f\in\arg\top_{k_{a}}\left(MS(\textbf{q},\textbf{f})\right)\right\},$ (8)
where $k_{a}$ is a hyperparameter that controls the number of raw mapped triple extracted, and f are obtained through encoding the text concatenation of entities and relation of the triplet fact.
Further, we apply LLM based filter on raw mapped triplet facts $\mathcal{F}_{\text{raw}}$ to generate the mapped triplet facts $\mathcal{F}$, to reduce the noise in mapped triples:
$\mathcal{F}=\text{LLM}(\mathcal{F}_{\text{raw}},q,p_{\text{tf}}),$ (9)
where $p_{\text{tf}}$ denotes the LLM prompt for triple filtering.
Next, based on the mapped triplet facts $\mathcal{F}$ and the KG structure, we calculate the Personalized PageRank (PPR) [33] score as our node influence score $\mathcal{S}_{\mathcal{V}}$ for node set $\mathcal{V}$:
$\mathcal{S}_{\mathcal{V}}=\lim_{i\to\infty}\text{PPR}^{i},\,\text{PPR}^{i}=(1-d)\times\mathbf{p}+d\times\mathbf{P}\cdot\text{PPR}^{i-1}.$ (10)
To align with [31], we set the damping factor $d$ to 0.5, and converge the PPR calculation when it changes less than $10^{-7}$; $\mathbf{p}\in\mathbb{R}^{|\mathcal{V}|}$ denotes the personalization vector that depends on the mapped triples $\mathcal{F}$. The $\mathbf{P}\in\mathbb{R}^{|\mathcal{V}\times\mathcal{V}|}$ in Equation 10 denotes the transition probability matrix, and can be obtained by:
$\mathbf{P}_{u,v}=\begin{cases}\frac{1}{|\mathcal{N}(u)|},&\text{if }e_{(u,v)}\in\mathcal{E},\\ \frac{1}{|\mathcal{V}|},&\text{otherwise},\end{cases}$ (11)
where $\mathcal{N}(u)$ denotes the 1-hop neighbor set of node $u$ in the KG. Moreover, following [31], after calculate the PPR score, we multiply the PPR score of all passage nodes by a balance factor 0.05, to avoid over concentration on them. After that, we can obtain the node influence score $s_{v}$ for each node $v\in\mathcal{V}$ as follow:
$s_{v}=\mathcal{S}_{\mathcal{V}}(v),\quad\forall v\in\mathcal{V}.$ (12)
Input :  $\mathcal{KG}(\mathcal{V},\mathcal{E})$: The indexed KG;
$\mathcal{F}$: A set of query-mapped triplet facts in $\mathcal{KG}$;
$\mathbf{q}$: The vector representation of the query $q$.
Output :  $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$: The weighted KG.
1 Assign node weights $\mathcal{S}_{\mathcal{V}}$ with PPR score w.r.t. $\mathcal{F}$ based on Equation 10: $\mathcal{S}_{\mathcal{V}}\leftarrow\lim_{i\to\infty}\text{PPR}^{i}$;
2 for $e_{(u,v)}\in\mathcal{E}$ do
3   Assign edge costs $c_{e_{(u,v)}}$ based on Equation 13: $c_{e_{(u,v)}}\leftarrow(1-\text{MS}(\mathbf{q},\mathbf{f}_{e_{(u,v)}}))/2$;
4 $\mathcal{C}_{\mathcal{E}}\leftarrow\mathcal{C}_{\mathcal{E}}\cup\{c_{e_{(u,v)}}\}$;
5
6return $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$;
Algorithm 2 Graph Weighting Algorithm
At last, we calculate the edge cost $\mathcal{C}_{\mathcal{E}}$ based on the similarity between query $q$’s representation q and relation edge $e$’s respective triplet facts $f_{e}$’s representation $\textbf{f}_{e}$:
$\mathcal{C}_{\mathcal{E}}=\{c_{e}\mid e\in\mathcal{E}\},\quad c_{e}=\frac{1-\text{MS}(\mathbf{q},\mathbf{f}_{e})}{2}.$ (13)
Moreover, to guarantee the KG’s connectivity, we connect all passage nodes to a pseudo node, with the pseudo edge (relation) namely pseudo_relation, and the cost of these pseudo edges is set to 10 to penalize paths through the pseudo node, aligning with the setting of HippoRAG2 [31].
Next, we introduce the Minimum Cost Maximum Influence (MCMI) subgraph generation problem, which can provide more comprehensive reasoning chains and result in more comprehensive answers in tasks that require knowledge summarization. According to Figure 3, the MCMI subgraph can include cyclic graph structures to cover key query related triples (colored in orange), which are not covered by simple node matching or Steiner tree, hence address the Poor Reasoning Ability and Inadequate Answer issue of current Graph-based RAG models. Here, we give a formal definition of the MCMI subgraph generation problem:
Definition 1 (Minimum Cost Maximum Influence Subgraph Generation Problem).
Given an node and edge weighted connected graph $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$, a set of mapped triples $\mathcal{F}$. The aim of this problem is to generate a subgraph $\mathcal{G}_{\text{MCMI}}$ with the maximum average node score and minimum total edge cost that contains all the mapped triples.
Next, we prove our MCMI Subgraph Generation problem can be reduced from the NP-hard Steiner Tree problem [70].
Theorem 1.
The Minimum Cost Maximum Influence Subgraph Generation problem is NP-hard.
Proof.
We define the decision version of MCMI as follows: Given a weighted knowledge graph $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$, entity node set from mapped triples (considered as terminal nodes) $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}$, and two thresholds $\alpha\in(0,1]$ and $\beta\in\mathbb{R}^{+}$, does there exist a connected subgraph $\mathcal{G}^{\prime}(\mathcal{V}^{\prime},\mathcal{E}^{\prime})\subseteq\mathcal{KG}$ such that:
  1. (i)
    $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}^{\prime}$, which means all mapped triples are included,
  2. (ii)
    the average node influence score satisfies $\frac{1}{|\mathcal{V}^{\prime}|}\sum_{v\in\mathcal{V}^{\prime}}s_{v}\geq\alpha$, and
  3. (iii)
    the total edge cost satisfies $\sum_{e\in\mathcal{E}^{\prime}}c_{e}\leq\beta$?
Now, consider an arbitrary MCST instance with graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, terminal node set $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}$, and total cost $\gamma$. We construct an MCMI instance as follows:
  1. 1.
    Use the same graph $\mathcal{G}$ as the KG $\mathcal{KG}$.
  2. 2.
    Set $s_{v}=1$ for all $v\in\mathcal{V}$, hence any subgraph has average node score exactly 1.
  3. 3.
    Keep edge costs $c_{e}$ unchanged for each $e\in\mathcal{E}$.
  4. 4.
    Set $\alpha=1$ and $\beta=\gamma$.
Because all nodes have score 1, condition (ii) holds for any non-empty subgraph. Therefore, a Steiner tree of cost $\leq\gamma$ exists in the MCST instance if and only if a feasible subgraph exists in the MCMI instance. This establishes a polynomial-time reduction, proving that MCMI is NP-hard. ∎
Since the MCMI Subgraph Generation problem is the NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, we propose a greedy algorithm (algorithm 3) based on the Mehlhorn’s algorithm [71] to solve this problem.
To begin with, we initialize the MCMI as the Minimum Cost Steiner Tree (MCST) through all mapped triplet facts $\mathcal{F}$ in the weighted KG $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$ (line 1 to line 10 of algorithm 3), where score $s_{v}\in\mathcal{S}_{\mathcal{V}}$ of each node $v$, and cost $c_{e}\in\mathcal{C}_{\mathcal{E}}$ of each edge $e$ are obtained with Equation 12 and Equation 13, respectively:
$\mathcal{G}_{\text{MCMI}}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}})=\text{MCST}(\mathcal{F},\mathcal{KG}),$ (14)
where $\text{MCST}(\mathcal{F},\mathcal{KG})$ denotes the Minimum Cost Steiner tree generation algorithm.
Next, we calculate the ratio of the nodes’ influence score and their connected edges’ cost (i.e., $s/c$), and check this ratio of all nodes in the current MCMI’s 1-hop neighborhood $\mathcal{N}(\mathcal{G}_{\text{MCMI}})$, if this ratio is greater than that the current MCMI subgraph’s average score-cost ratio, we add this node (entity) and the corresponding edge (relation) to the MCMI, recalculate the MCMI’s score-cost ratio and continue the iteration, until all neighbor node’s score-ratio is less or equal than the MCMI subgraph’s average score-cost ratio. That is, for each edge $e_{(u,v)}$, we have (line 10 to line 16 of algorithm 3):
[…]
where $r_{\text{MCMI}}$ denotes the average score-cost ratio of the current MCMI subgraph; $\mathcal{V}_{\text{MCMI}}$ and $\mathcal{E}_{\text{MCMI}}$ denotes the node and edge set of current MCMI subgraph; $e_{(\text{MCMI},v)}$ denotes the outgoing edges from the MCMI subgraph to its 1-hop neighborhood; And $r_{\text{MCMI}}$ is calculated as follow:
$r_{\text{MCMI}}=\sum_{e_{(u,v)}\in\mathcal{E}_{\text{MCMI}}}\frac{s_{u}+s_{v}}{2\times c_{e_{(u,v)}}\times|\mathcal{E}_{\text{MCMI}}|}.$ (16)
After generate MCMI subgraph generation, we serialize the generated MCMI subgraph into a text graph string $g$:
$g=\text{StrConcat}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}}).$ (17)
Running Example: As shown in Figure 3, if we simply calculate the minimum cost Steiner tree within the given KG, the algorithm will simply return the two mapped fact (Kidney, Work, Kidney cancer), and (Kidney, Spread, Tumors) as the reasoning path, and the LLM’s response answer would likely to be Tumors. However, if we go one step further, then within the generated cyclic graph, we can have more complete reasoning information regarding the query, and involve the correct answer Blood vessels.
Theorem 2.
The approximation ratio of the MCMI Subgraph Generation Algorithm (algorithm 3) is 2.
Proof.
In algorithm 3, we firstly compute the minimum cost Steiner tree $\mathcal{G}(\mathcal{V}_{\text{MCST}},\mathcal{E}_{\text{MCST}})$ which connects all nodes in $\mathcal{V}_{\text{term}}$. Suppose the total cost for $\mathcal{G}$ is $\gamma$, and the optimal minimum cost Steiner tree is $\mathcal{G}^{*}(\mathcal{V}_{\text{MCST}^{*}},\mathcal{E}_{\text{MCST}^{*}})$ with cost $\gamma^{*}$. $\mathcal{G}^{*}$ covers at least all edges from $\mathcal{G}$. Following triangle inequality, for $\forall e\in\mathcal{E}_{\text{MCST}^{*}}$, the sum of weight is less than 2 times of edge weights in $\mathcal{E}_{\text{MCST}}$, which is $\gamma\leq 2\gamma^{*}$. The process from line 12 to line 15 will retrieve more edges. However, they will not change the weight in $\mathcal{G}$. Hence, the approximation ratio of algorithm 3 is 2. ∎
Input :  $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$: The weighted KG;
$\mathcal{F}$: A set of query-mapped triplet facts in $\mathcal{KG}$.
Output :  $\mathcal{G}_{\text{MCMI}}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}})$: The MCMI subgraph.
1 Compute a minimum cost Steiner tree MCST $\mathcal{G}$ of $\mathcal{KG}$;
2 Let $\mathcal{V}_{\text{term}}\leftarrow\mathcal{F}\cap\mathcal{V}_{\text{MCST}}$ be the terminal nodes in MCST;
3 Construct weighted auxiliary graph $H$: $\mathcal{V}_{H}\leftarrow\mathcal{V}_{\text{term}}$;
4 for each pair of terminal nodes $u,v\in\mathcal{V}_{\text{term}}$ do
5   Find the shortest path $P_{(u,v)}$ in $\mathcal{G}$ from $u$ to $v$;
6   Let $c^{H}_{(u,v)}\leftarrow\sum_{e\in P_{(u,v)}}c_{e}$, where $c_{e}\in\mathcal{C}_{\mathcal{E}}$;
7
8Compute minimum cost Steiner tree $\mathcal{G}^{\prime}$ of $H$;
9 Let $\mathcal{V}_{\text{MCMI}}\leftarrow\emptyset$, $\mathcal{E}_{\text{MCMI}}\leftarrow\emptyset$;
10 for each edge $e_{(u,v)}\in\mathcal{G}^{\prime}$ do
11   Add nodes and edges from shortest path $P_{(u,v)}$ in MCST to $\mathcal{V}_{\text{MCMI}}$ and $\mathcal{E}_{\text{MCMI}}$;
12
13Calculate $r_{\text{MCMI}}$ with Equation 16;
14 while the node $v$ in MCMI’s 1-hop neighborhood with highest $s_{v}/c_{e_{(\text{MCMI},v)}}$ satisfies $s_{v}/c_{e_{(\text{MCMI},v)}}>r_{\text{MCMI}}$ do
15   Recalculate $r_{\text{MCMI}}$ with Equation 16;
16 $\mathcal{V}_{\text{MCMI}}\leftarrow\mathcal{V}_{\text{MCMI}}\cup\{v\}$;
17 $\mathcal{E}_{\text{MCMI}}\leftarrow\mathcal{E}_{\text{MCMI}}\cup\{e_{(\text{MCMI},v)}\}$;
18
19return $\mathcal{G}_{\text{MCMI}}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}})$;
Algorithm 3 MCMI Subgraph Generation Algorithm
III-D Step 3: Hybrid Retrieval
Other than Graph retrieval, we also apply Hybrid Retrieval (HR) to further enrich the context for query answering. In Hybrid Retrieval, we retrieve the most similar text chunks $\mathcal{T}_{HR}$ from all text chunks $t\in\mathcal{T}_{c}$ to the query’s vector representation q, based on the hybrid retrieval score $\text{HS}(\cdot)$:
$\mathcal{T}_{\text{HR}}=\left\{t\in\mathcal{T}_{c}\,\middle|\,f\in\arg\top_{k_{r}}\left(\text{HS}({q},{t})\right)\right\},$ (18)
$\text{HS}({q},{t})=\frac{\text{MS}(\textbf{q},\textbf{t})+\text{BM25}(q,t)}{2}$ (19)
where $k_{r}$ denotes the hyperparameter that controls the number of hybrid retrieved text chunks, t, q denotes the vector representation of text chunk $t$ and query $q$, respectively; $\text{HS}(q,t)$, $\text{BM25}(q,t)$ and $\text{MS}(\textbf{q},\textbf{t})$ denotes the hybrid retrieval score, the BM25 [38] sparse retrieval score, and the dense representation vectors’ cosine similarity score between query $q$ and text $t$ in Equation 7.
After obtaining the textual MCMI subgraph string $g$ and the chunks $\mathcal{T}_{\text{HR}}$ from hybrid retrieval, we input them together with the query $q$ and the the query answer generation LLM instruction $p_{\text{gen}}$, to ask the LLM to give an answer to query $q$ based on these provided contents.
$\text{answer}=\text{LLM}(q,g,\mathcal{T}_{\text{HR}},p_{\text{gen}}),$ (20)
III-E Complexity Analysis
III-E1 Time Complexity
For the entity extraction procedure, the time complexity would be $O(|\mathcal{V}||\mathcal{T}_{c}|)$; Adding passage nodes also requires $O(|\mathcal{V}||\mathcal{T}_{c}|)$ time. For graph weighting, calculate the PPR score requires $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|))$ time, and calculate the edge cost requires $O(|\mathcal{E}|)$ time. For AGRAG’s Step 3, the time complexity of AGRAG’s MCST generation is $O(|\mathcal{E}|+|\mathcal{V}|log|\mathcal{V}|)$ using the Mehlhorn algorithm [71]; The expansion from MCST to MCMI would cost $O(|\mathcal{V}_{\text{MCMI}}|\cdot|\mathcal{N}(\mathcal{V}_{\text{MCMI}})|)$ Hence, the time complexity of AGRAG’s MCMI generation at the worst case when $\mathcal{V}_{\text{MCMI}}=\mathcal{V}$ and the graph is fully connected would be: $O(|\mathcal{V}|^{2}+|\mathcal{E}|).$
Hence, the total time complexity of AGRAG is: $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|)+|\mathcal{V}|^{2})$. Since the number of text chunk $|\mathcal{T}_{c}|$ is less or equal to the number of nodes $|\mathcal{V}|$ in graph, in the worst case, the total time complexity of AGRAG is:
$O(|\mathcal{V}|\cdot|\mathcal{E}|+|\mathcal{V}|^{2})$.
III-E2 Space Complexity
The graph of AGRAG’s Step 1 and Step 2 is stored with the dense weighted adjency matrix, which needs $O(|\mathcal{E}|)$ space; Storing the text chunk would need $O(|\mathcal{T}_{c}|)$ space; Storing the entity nodes require $O(|\mathcal{V}|)$ space. Hence the total space complexity of AGRAG’s Graph Indexing step would be: $O(|\mathcal{V}|+|\mathcal{E}|+|\mathcal{T}_{c}|)$. In AGRAG’s Step 3, we only need $O(|\mathcal{E}|)$ to store the graph.
Hence, the total space complexity of AGRAG is:
$O(|\mathcal{V}|+|\mathcal{E}|+|\mathcal{T}_{c}|)$.
IV Experiments
In this section, we present the experimental evaluation of our framework AGRAG against 6 effective baselines spanning 3 technical categories across 5 tasks. Specifically, we first outline the experimental setup in Section IV-A, including details on datasets, and evaluation metrics. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section IV-B and Section IV-C, respectively. Finally, we conduct ablation study in subsection IV-D, parameter sensitivity study in subsection IV-E, and case study in subsection IV-D.
IV-A Experiment Settings
IV-A1 Datasets & Tasks
We select the Novel and Medical dataset from GraphRAG-bench benchmark [34] to evaluate AGRAG’s performance on 4 tasks that have different complexity levels, these tasks have the same input and output format with the classic QA task:
  • Fact Retrieval: The simplest (level 1) task of GraphRAG-bench, to solve this task, the model needs to retrieve isolated facts in the graph with minimal reasoning needed. This task mainly tests the models’ precise keyword matching ability.
  • Complex Reasoning (Complex Rea.): The level 2 task of GraphRAG-bench, to solve this task, the model needs to connect multiple facts on the graph to form multi-hop, long-ranged reasoning chains.
  • Contextual Summarization (Contextual Sum.): The level 3 task of GraphRAG-bench. Here, the model needs to synthesize fragmented information from the graph into comprehensive summaries.
  • Creative Generation (Creative Gen.): The most complex (level 4) task of GraphRAG-bench, to solve this task, the model needs to generate hypothetical responses based on but beyond the retrieved knowledge.
Furthermore, we also apply the text classification task, to test the RAG models’ ability to summarize the topic of text passages, and classify them into the correct category within a given label sets.
  • Text Classification: Give a set of texts $\mathcal{T}$ and a set of label $\mathcal{L}$, the classification task aim to assign each text $t\in\mathcal{T}$ a label $l\in\mathcal{L}$ that best suited to describe its semantic meaning. For text classification task, we select the Web-Of-Science (WOS) dataset [72], and the IFS-Rel dataset [73].
We select one text passage per label from the training sets of the WOS and IFS-Rel datasets to construct the corpus for RAG. During retrieval, instead of retrieving the full text chunk, we enable HippoRAG2 and AGRAG to directly retrieve the corresponding label of each text chunk. The retrieved labels are then used to form a candidate label set for the LLM’s text prediction. The LLM is subsequently tasked with selecting the most appropriate label from this candidate set.
The detailed dataset and task statistics are shown in Table II.
TABLE II: Statistics of the Novel and Medical dataset from GraphRAG-bench [34] across 4 tasks, and statistics of the WOS [72] and IFS-REL [73] dataset for the text classification task.
Dataset Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval
# of Questions
Novel 67 362 289 971
Medical 166 289 509 1,098
Dataset Text Classification
# of Texts # of Labels Token Per Text
WOS 9,396 133 200
IFS-REL 2,560 64 105
TABLE III: Model effectiveness on 4 datasets for 5 tasks, all experiments were carried out with Qwen2.5-14B-Instruct LLM backbone.
Category Task Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval Text Classification
Dataset Novel Medical Novel Medical Novel Medical Novel Medical WOS IFS-REL
Metric FS. ACC. COV. FS. ACC. COV. COV. ACC. COV. ACC. ROG. ACC. ROG. ACC. ROG. ACC. ROG. ACC. ACC. REC. ACC. REC.
Traditional
RAG
NaiveRAG 0.205 0.307 0.384 0.176 0.577 0.554 0.745 0.621 0.677 0.576 0.120 0.447 0.117 0.635 0.117 0.431 0.182 0.604 0.237 0.257 0.351 0.197
L.LLMLingua 0.213 0.316 0.377 0.225 0.575 0.496 0.687 0.627 0.657 0.583 0.092 0.434 0.080 0.633 0.115 0.420 0.105 0.591 0.274 0.231 0.378 0.220
Community
-based
GraphRAG 0.427 0.310 0.339 0.345 0.475 0.325 0.758 0.620 0.625 0.506 0.246 0.449 0.250 0.606 0.240 0.445 0.217 0.521 0.253 0.201 0.279 0.202
LightRAG 0.201 0.314 0.329 0.189 0.567 0.549 0.753 0.628 0.625 0.585 0.223 0.448 0.255 0.637 0.253 0.460 0.283 0.607 0.273 0.224 0.265 0.220
Graph-
wandering
HippoRAG2 0.496 0.295 0.325 0.347 0.563 0.451 0.735 0.586 0.496 0.560 0.314 0.453 0.332 0.622 0.310 0.470 0.343 0.598 0.424 0.270 0.389 0.246
F.GraphRAG 0.286 0.287 0.376 0.217 0.500 0.396 0.706 0.528 0.507 0.587 0.237 0.371 0.262 0.643 0.318 0.439 0.352 0.610 0.283 0.245 0.307 0.221
AGRAG 0.513 0.339 0.386 0.356 0.598 0.589 0.778 0.643 0.690 0.604 0.319 0.467 0.343 0.665 0.324 0.490 0.357 0.628 0.438 0.307 0.403 0.251
TABLE IV: The performance comparasion of AGRAG and baselines with GPT-4o [74] and R1-Searcher-Qwen-2.5-7B [75] backbone
Backbone Task Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval
Dataset Novel Medical Novel Medical Novel Medical Novel Medical
Metric FS. ACC. COV. FS. ACC. COV. COV. ACC. COV. ACC. ROG. ACC. ROG. ACC. ROG. ACC. ROG. ACC.
GPT-4o HippoRAG2 0.519 0.429 0.355 0.322 0.571 0.545 0.752 0.612 0.671 0.597 0.340 0.477 0.391 0.667 0.340 0.490 0.400 0.631
AGRAG 0.536 0.478 0.414 0.374 0.632 0.609 0.794 0.667 0.703 0.634 0.351 0.515 0.397 0.673 0.343 0.503 0.402 0.649
R1-Searcher NaiveRAG 0.211 0.203 0.255 0.210 0.036 0.124 0.604 0.444 0.181 0.254 0.069 0.285 0.050 0.200 0.086 0.306 0.063 0.208
HippoRAG2 0.336 0.187 0.259 0.275 0.258 0.250 0.708 0.483 0.456 0.416 0.184 0.304 0.231 0.448 0.215 0.272 0.242 0.490
AGRAG 0.400 0.255 0.263 0.308 0.270 0.261 0.715 0.495 0.480 0.429 0.208 0.326 0.245 0.464 0.313 0.303 0.253 0.511
IV-A2 Evaluation Metrics
In this paper, we apply 5 metrics to evaluate AGRAG’s performance on 5 different tasks:
  1. a.
    Accuracy (ACC.): For text classification, the proportion of exact matches between LLM and gold labels. For other tasks, sample is scored 0, 0.5, or 1 by the LLM for inaccurate, partially accurate, or exactly accurate responses.
  2. b.
    Recall (REC.): Quantifies the model’s ability to identify all actual positive instances by calculating the ratio of true positives to the sum of true positives and false negatives, we calculate the macro recall score and apply it to the text classification task, where we have a fixed label set.
  3. c.
    ROUGE-L (ROG.) [76]: Evaluates the answer quality of by comparing the longest common subsequence (LCS) between the generated answer $G$ and the gold answer $Y$.
    $\text{ROUGE-L}=\frac{\text{LCS}(G,Y)}{|G|},$ (21)
    we apply ROUGE-L to the fact retrieval and the complex reasoning tasks, as the query answer of these two tasks are usually a fixed entity or relation.
  4. d.
    Coverage (COV.) [34]: Evaluates answer completeness by assessing whether all necessary reference evidences $E$ are included in the LLM’s response $R$, computed as the proportion of covered evidences to the number of required evidences $|E|$:
    $\text{Cov}=\frac{|\{e\in E\ |\ M(e,R)\}|}{|E|},$ (22)
    where $M(e,R)$ is a boolean function indicating whether evidence $e$ is covered by generated response $R$. We apply the coverage metric to contextual summarization and creative generation tasks, as the their answers are usually passages that conclude multiple information points.
  5. e.
    Faithfulness (FS.) [34]: Measures the percentage of answer claims $A$ directly supported by retrieved context $C$, calculated as the ratio of claims aligned with evidence to the number of total claims $A$:
    $\text{FS}=\frac{|\{a\in A\ |\ S(a,C)\}|}{|A|},$ (23)
    where $S(a,C)$ is a boolean function indicating whether a claim is supported by contexts in $C$. We apply the faithfulness metric to the creative generation task to evaluate the quality of the hypothesis in the LLM’s response.
TABLE V: Model efficiency on 4 datasets for 5 tasks, all experiments were carried out with Qwen2.5-14B-Instruct LLM backbone. Where AGRAG achieve at most 166.3% accleration, and 368.8% less token cost compared with tested baselines.
Category Models Time Cost (mins) Token Cost (# per query)
Novel Medical WOS IFS-REL Novel Medical WOS IFS-REL
Traditional
RAG
NaiveRAG 384.5 (+85%) 252.3  (+63%) 294.2 (+98%) 227.1 (+64%) 11,145 (+46%) 7,852 (+41%) 7,321 (+80%) 4,537 (+129%)
L.LLMLingua 410.5 (+98%) 342.8 (+122%) 362.2 (+144%) 318.7 (+13%) 10,216 (+34%) 6,575 (+18%) 5,892 (+45%) 2,771 (+40%)
Community
-based
GraphRAG 552.6 (+166%) 352.1 (+128%) 325.0 (+119%) 281.7 (+104%) 14,795 (+94%) 11,511 (+107%) 12,160 (+199%) 9,296 (+369%)
LightRAG 386.5 (+86%) 279.5 (+81%) 252.7 (+70%) 229.4 (+66%) 14,377 (+89%) 9,961 (+79%) 11,737 (+188%) 7,866 (+297%)
Graph-
wandering
HippoRAG2 229.2 (+11%) 176.1 (+14%) 190.4 (+28%) 154.7 (+12%) 8,978 (+18%) 6,859 (+23%) 5,159 (+27%) 2,542 (+28%)
F.GraphRAG 319.8 (+54%) 221.44 (+43%) 215.4 (+45%) 191.1 (+38%) 8,742 (+15%) 6,076 (+9%) 4,992 (+23%) 2,237 (+13%)
AGRAG 207.5 154.4 148.6 138.4 7,627 5,558 4,071 1,983
IV-A3 Baselines
We compare AGRAG’s performance with 6 baselines from 3 technical categories, they are traditional RAG model NaiveRAG [15] and LongLLMLingua [69]; Community-based model GraphRAG [25], and LightRAG [26]; And walking-based model Fast-GraphRAG [32] and HippoRAG2 [31]. We use the Qwen-2.5-14B-Instruct [77] as the LLM backbone unless further illustrated.
Traditional RAG Models These models directly retrieve text chunks based on the semantic similarity between chunks and the query.
  • NaiveRAG [15]: NaiveRAG acts as a foundational baseline of current RAG models. It encodes text chunks into a vector database and retrieve the Top-k text chunks based on the vector similarities between queries and text chunks.
  • L.LLMLingua [69]: LongLLMLingua utilizes LLM’s generation perplexity to filter out unimportant LLM input tokens. In this paper, we apply it to compress the retrieved text chunks of NaiveRAG w.r.t. the query and task instruction.
Community-based Models These models retrieve one or multi-hop neighborhood of certain graph nodes that correspond to the query.
  • GraphRAG [25]: It employs LLM for graph construction. When retrieval, it aggregates nodes into communities w.r.t. the query, and generates community reports to encapsulate global information from texts.
  • LightRAG [26]: It calculates the vector similarity between query extracted entities and graph nodes, achieving a one-to-one mapping from entities to graph nodes, then retrieve these nodes’ graph neighborhood.
Walking-based Models These models apply walking-based graph algorithms to score graph nodes, then retrieve text chunks w.r.t. their contained entity nodes’ score.
  • F.GraphRAG [32]: Fast-Graphrag employs LLM for graph construction. During retrieval, it applies PPR to assign node scores based on the query, and retrieve the Top-k passages based on the sum of their contained nodes’ scores.
  • HippoRAG2 [31]: It employs LLM for graph construction, and also add the text chunks as nodes in the graph. As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.
IV-A4 Hyperparameter and Hardware Settings
We set the retrieval top-k value for all RAG models to 5, aligning with the settings of Fast-GraphRAG and HippoRAG2. For GraphRAG, we apply its local search mode to all tasks, as it achieves the overall best effectiveness. For LightRAG, we apply its hybrid search mode for all tasks except the text classification task, as it is claimed as one of the contributions of LightRAG. For Fast-GraphRAG and HippoRAG2, we set their hyperparameters to default, to align with their original paper and code. Other hyperparameters of models were set as default by their codes.
For our AGRAG, the length of text chunks $l_{t}$ is set to 256, and token overlap between chunks $l_{o}$ to 32, the number of raw mapped triples $k_{a}$ and the number of hybrid retrieved chunks $k_{r}$ are both set to 5, to align with the setting of HippoRAG2 and GraphRAG-bench. AGRAG’s unique hyperparameter, the entity extraction threshold, and the maximum n-gram, were set to 0.5 and 3, respectively. Please refer to subsection IV-E for further parameter sensitivity analysis.
All our experiments were carried out and evaluated with Qwen2.5-14B-Instruct [77] LLM backbone if not further illustrated. We apply the Contriever [78] embedding model for HippoRAG2 and AGRAG’s dense passage retrieval procedure, and bge-large-en-v1.5 [79] embedding model for other baselines’ LLM embedding. All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-80GB GPU.
IV-B Effectiveness Evaluation
Other than Qwen-2.5-14B-Instruct, we also test the performance of AGRAG with GPT-4o [74] and the reinforcement learning based RAG reasoning backbone R1-Searcher-Qwen-2.5-7B [75] in Table IV. Due to cost and space considerations, we only compare AGRAG’s performance with the previous state-of-the-art model HippoRAG2 [31] on GraphRAG-bench benchmark [34] for GPT-4o, as the text classification task has a large number of texts to be classified, and can lead to high token cost (Over 7,000 USD on WOS dataset, \$15 for 1 million tokens). For R1-Searcher, we further compare NaiveRAG [15], as it is the RAG framework used in its original paper [75]. Based on the experiment results in Table III and Table IV, we have observations as follow:
Firstly, in line with recent findings [34, 35, 36], NaiveRAG [15] achieves better performance than all tested walking-based baselines on Contextual Summarization and Creative Generation tasks. This is because, walking-based baselines overlook the global, and coarse-grained information that is essential for effective summarization and creative generation. This result also suggests that the fine-grained, phrase-level information they capture is less beneficial for summarization compared to higher-level, coarse-grained insights, which help the LLM form a holistic understanding of the text chunks. Our AGRAG achieves the best performance on these two tasks by providing the reasoning subgrah, as well as the hybrid retrieved text chunks to the LLM, thereby offering more comprehensive coarse to fine-grained context for query answering.
Secondly, on simpler Complex Reasoning and Fact Retrieval tasks, the more recent walking-based models achieve the best performance. This indicates that the phrase-level dependencies, both long and short range, captured by these models’ graph scoring mechanisms, are beneficial for solving tasks that require forming a reasoning chain from one or multiple facts to provide factual information. Our AGRAG achieves the best performance on Fact Retrieval and Complex Reasoning tasks. as it can benefit from phrase-level dependencies by generating explicit reasoning chains and directly feeding it to the LLM as guidance.
Lastly, in the text classification task, HippoRAG2 [31] achieves the second-best performance across all metrics, slightly outperformed only by AGRAG. This is because text classification requires capturing correlations between specific text entities and the text chunks to be classified, and both HippoRAG2 and AGRAG explicitly model entity-to-passage relationships in the KG. AGRAG achieves the best performance as its reasoning subgraph spans multiple entity nodes in the KG and explicitly captures the relationships between these entities and the text chunks associated with a given classification label, which can assist LLM classification.
IV-C Efficiency Evaluation
TABLE VI: The ablation experiment result on the two datasets and 4 tasks from GraphRAG-bench [34] benchmark. All experiments were carried out with Qwen2.5-14B-Instruct LLM backbone.
Model Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval Total Efficiency
Novel Medical Novel Medical Novel Medical Novel Medical Novel Medical
FS. COV. FS. COV. COV. ROUGE-L Time Token Time Token
AGRAG w. LLM ER 0.4275 0.3650 0.3519 0.5873 0.7553 0.6000 0.3196 0.3330 0.3220 0.3525 244.50 8,917 180.71 5,971
AGRAG w.o. HR 0.5194 0.3433 0.3419 0.5701 0.6771 0.4199 0.3175 0.3240 0.3136 0.3435 200.64 5,615 120.79 5,135
AGRAG w.o. MCMI 0.4882 0.3580 0.3506 0.5749 0.7479 0.6003 0.3242 0.3292 0.3221 0.3587 194.27 7,467 134.50 5,561
AGRAG 0.5134 0.3857 0.3560 0.5888 0.7779 0.6904 0.3186 0.3431 0.3241 0.3572 207.45 7,627 154.42 5,558
To evaluate the efficiency of AGRAG, we compare AGRAG’s total time and token cost with other baselines’ on all 4 dataset tested, the results are shown in Table V. According to the result, AGRAG can also achieve the best time and token efficiency compared with other baselines. It achieves at most 1.66 times faster compared with GraphRAG on Novel dataset, and at least 10.5% accleration compared to HippoRAG2 on Novel dataset. As for token saving, AGRAG also achieves at most 3.69 times token saving compared with GraphRAG on IFS-REL dataset, and save 9.3% tokens at least compared with the previous state-of-the-art model HippoRAG2.
These results lead to two major findings: Firstly, both HippoRAG2 and AGRAG achieve superior time efficiency than other baselines, this suggests that frequent LLM calls, particularly those involving long input sequences, are the primary efficiency bottleneck in most Graph-based RAG models. By avoiding the LLM-based graph summarization process, HippoRAG2 and AGRAG significantly alleviate this bottleneck, leading to improved performance.
Secondly, although AGRAG introduces an additional procedure generating the MCMI reasoning graph, the statistics-based entity extraction method avoids the LLM calling. Denote the number of nodes in graph as $|\mathcal{V}|$, the number of nodes in text chunk $t$ as $|\mathcal{V}_{t}|$, the token length of text chunk $t$ as $l_{t}$, the dimension of LLM hidden state is $d_{\text{llm}}$, for each test chunk, AGRAG’s statistics-based entity extraction achieves $O(|\mathcal{V}|)$ time complexity, and the LLM entity extraction method achieves $O(d_{\text{llm}}(|\mathcal{V}_{t}|+l_{t})^{2})$ complexity. As $d_{\text{llm}}$ is very large for LLM models (e.g., over 12,000 for GPT-4o [74]), and $|\mathcal{V}_{t}|+l_{t}$ can also be a large number, our AGRAG can still achieve better overall efficiency.
IV-D Ablation Study
To analyze AGRAG’s performance, we conduct an extensive ablation study on GraphRAG-bench tested across 2 datasets and 4 tasks, the ablation models tested is listed as follow:
  • AGRAG w. LLM ER: Here, we substitute AGRAG’s statistics-based entity extraction method for HippoRAG2’s LLM entity extraction approach to test its effects.
  • AGRAG w.o. HR: Here, we remove the hybrid retrieved chunks in AGRAG’s LLM input, and only input the reasoning subgraph and its involved text chunk to LLM.
  • AGRAG w.o. MCMI: Here, we simply generate a minimum cost Steiner tree that contains all mapped fact triples as reasoning path, ignoring more complex reasoning structures.
Based on the results in Table VI, we can have the following observations regarding AGRAG:
Firstly, although AGRAG w. LLM ER is more time and token-consuming, it still achieves worse performance compared to AGRAG. Based on the case study in subsection IV-F. This is because the LLM entity extraction approach can produce hallucinated outputs that fail to conform to the expected entity extraction format, leading to decoding errors. In this scenario, current models either skip these entities or text chunks, resulting in incomplete knowledge, or construct graph with the hallucinated outputs, leading to low graph construction quality. This highlights the importance of substituting the LLM entity extraction method to avoid hallucination.
(a) Parameter sensitivity experiment on Novel dataset of GraphRAG-bench.
(b) Parameter sensitivity experiment on Medical dataset of GraphRAG-bench.
Figure 4: Parameter sensitivity experiment on GraphRAG-bench, where x-axis of each heatmap denotes the entity extraction threshold, and y-axis denotes the maximum n-gram. We choose accuracy as the metric; the darker the color of each block, the better the accuracy of the corresponding parameter pair.
Secondly, AGRAG w.o. HR experiences a significant performance drop on the Contextual Summarization task and in the coverage score of the Creative Generation task. This is because it provides the LLM with only the reasoning path and its associated text chunks—while efficient, this limits the amount of supportive content available and restricts the model’s ability to find sufficient grounded information.
At last, AGRAG w.o. MCMI performs similarly to AGRAG on the Fact Retrieval and Complex Reasoning tasks. This is because its simple, tree-structured reasoning paths are insufficient to support queries from harder tasks, which may require the inclusion of cyclic structures or multiple answer nodes that are difficult to jointly capture within the Steiner tree framework.
IV-E Parameter Sensitivity Analysis
(a) A comparasion between the LLM entity extraction of HippoRAG2 and AGRAG’s statistic-based entity extraction. Compared with LLM entity extraction method, our method extracts more candidate entities, lead to generating more natural and semantically complete triples.
(b) A comparison between generating the Steiner tree and the MCMI for reasoning, where the red edges denote those relations within both the Steiner tree and the MCMI, and the orange edges denote relations solely within the MCMI. In this case, the Steiner tree does not retrieve the usage example and fail to answer the second question.
Figure 5: Two cases of AGRAG w.r.t. its statistics-based entity extraction method and MCMI generation step from GraphRAG-bench’s novel dataset.
In this subsection, we analyse the parameter sensitivity of AGRAG’s only two unique hyperparameters: The maximum n-gram $b$ and the entity extraction threshold $\tau$. As all other hyperparameters of AGRAG also exist in other baselines, we just align them with other baselines for fair comparison. We carry out the analysis on GraphRAG-bench [34], which cover most of the tasks of our experiment.
IV-E1 The maximum n-gram $b$
As introduced in subsection III-B, the maximum n-gram $b$ is used as the entity boundary to control the maximum allowed word length of the extracted entities. Here, we grow a test about the maximum n-gram parameter $b\in\{1,3,5,7\}$, as entity distribution in English is, in general, less than 7 words according to [80]. From the experiment results in Figure 4, AGRAG achieves the overall best performance when the maximum n-gram $b$ is set to 3, this is because most of the entities usually consist of no more than 3 words, according to the discovery of [81, 82, 80].
For Creative Generation task, AGRAG can achieve better performance with larger maximum n-gram, because the correct answer of this task are more flexiable than other tasks that have fixed answers or require answers to cover certain facts. Therefore, the negative impact of noise introduced by high n-gram phrases is overwhelmed by the additional context they provide, leading to better performance. This also highlights AGRAG’s adaptability to different tasks through its flexible graph construction. In our experiments, we set the maximum n-gram $b$ to 3 for all tasks.
IV-E2 The entity extraction threshold $\tau$
As introduced in subsection III-B, in the extraction procedure, we extract words or phrases that have the entity extraction score $\text{ER}(\cdot)>\tau$ as an entity, and we have $0<\text{ER}(\cdot)<1$. Considering the computational cost, we only test the entity extraction threshold $\tau\in\{0.1,0.3,0.5,0.7\}$. According to the experiment results in Figure 4, AGRAG typically achieves its best performance when set to 0.3 or 0.5. This is because a large threshold tends to extract less informative entities, and a small one tends to cause insufficient entity extraction; both can affect the model’s effectiveness. In our experiments, we set the entity extraction threshold $\tau$ to 0.5 for all tasks.
IV-F Case Study
In this subsection, we examine explicit case studies from our testing dataset. Due to space constraints, we focus on example cases related to AGRAG’s statistics-based entity extraction and its MCMI reasoning subgraph generation approach.
Firstly, we compare the entity extraction and triple extraction results of AGRAG with those of HippoRAG2, which employs an LLM-based entity extraction method but shares the same triple extraction process as AGRAG. As shown in Figure 5(a), the statistics-based entity extraction approach can avoid entity extraction failures caused by LLM hallucinations, and extract more candidate entities from the given text chunk, leading to the extraction of more natural and semantically complete triples.
Secondly, we compare AGRAG’s MCMI subgraph with a simple Steiner tree. In Figure 5(b), the Steiner tree is represented by red edges, while the MCMI subgraph includes both red and orange edges. Both reasoning structures successfully answer the first question marked in red in the query. However, the Steiner tree fails to address the second question marked in orange. In contrast, the MCMI subgraph incorporates examples of symbolic usage, such as predilection number for Mayas, enabling the LLM to generate more comprehensive and nuanced answers.
V Conclusion
This paper presents AGRAG, an Advanced Graph-based Retrieval Augmented Generation framework addressing the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current RAG frameworks. AGRAG introduces a statistics-based entity extraction method to replace the LLM entity extraction method applied by current baselines, avoiding hallucination, improving effectiveness, and having better efficiency. By formulating retrieval as a Minimum Cost Maximum Influence (MCMI) subgraph generation problem, AGRAG captures further query-relevant semantic dependencies by allowing more complex graph structure, and the reasoning paths can be more comprehensive and include more query-related graph nodes. By providing explicit reasoning paths to help LLM inference, AGRAG makes the LLM better focus on query-related contexts within the retrieved context. These can lead to better performance on complex tasks that require content summarization. Experiments on real-world datasets demonstrate the effectveness and efficiency of AGRAG over state-of-the-art RAG models.
VI AI-Generated Content Acknowledgement
In this paper, we use AI models, including but not limited to Qwen3-Max and Claude-4.5-sonnet, to polish writing as well as code debugging.